OpenStudy (anonymous):

A pebble is dropped into a graduated cylinder and the volume rises from 25.0 mL to 31.0 mL. The mass of the pebble is 3.5g. What is the density of the pebble?

OpenStudy (anonymous):

the difference is 6

OpenStudy (anonymous):

d=m/v so you do 3.5 over what

OpenStudy (anonymous):

thanks @abb0t ......

OpenStudy (abb0t):

Ok, so density = \(\frac{mass}{volume}\)

OpenStudy (abb0t):

But notice that you have two volumes. So that means, find the difference in the volume (meaning to subtract). That will give you the approximate volume of the pebble to use to find density.

OpenStudy (abb0t):

Does that make sense? And volume cannot be negative! Reemember that. So you are taking the bigger volume and subtracting the lesser volume.

OpenStudy (anonymous):

once you find the difference how do you use it in the equation?

OpenStudy (anonymous):

the difference is 6 so do i do 3.5/6 ?

OpenStudy (abb0t):

I will use \(D\) to represent density. D = \(\frac{3.5}{31.0-25.0}\)

OpenStudy (abb0t):

Yes

OpenStudy (anonymous):

so its 0.583

OpenStudy (abb0t):

That is your final answer.

OpenStudy (anonymous):

Thank you @abb0t !!!!!!